### v1

# Open Vocabulary Monocular 3D Object Detection

Abstract
In this work, we pioneer the study of open-vocabulary monocular 3D object detection, a novel task that aims to detect and localize objects in 3D space from a single RGB image without limiting detection to a predefined set of categories. We formalize this problem, establish baseline methods, and introduce a class-agnostic approach that leverages open-vocabulary 2D detectors and lifts 2D bounding boxes into 3D space. Our approach decouples the recognition and localization of objects in 2D from the task of estimating 3D bounding boxes, enabling generalization across unseen categories. Additionally, we propose a target-aware evaluation protocol to address inconsistencies in existing datasets, improving the reliability of model performance assessment. Extensive experiments on the Omni3D dataset demonstrate the effectiveness of the proposed method in zero-shot 3D detection for novel object categories, validating its robust generalization capabilities. Our method and evaluation protocols contribute towards the development of open-vocabulary object detection models that can effectively operate in real-world, category-diverse environments.
1 Introduction
Recognizing objects from a single image has been a longstanding task in computer vision, with broad applications in robotics and AR/VR. Over recent decades, 2D object detection — identifying and localizing objects within a 2D image plane — has achieved substantial progress, driven by advancements in deep learning techniques [16, 44, 20, 9] and large annotated datasets [31, 19, 49]. However, recognizing only a fixed set of objects is limiting, given the vast diversity of objects in real-world settings; detecting objects solely in 2D space is also insufficient for most practical tasks, as the world and its objects exist in 3D space.
To address these challenges, one line of recent research has focused on open-vocabulary 2D object detection [32, 12, 55, 65] (Fig. 1b) to identify objects beyond a fixed set of categories. Another line explores the monocular 3D detection task [5, 47, 53, 37] (Fig. 1c), which extends detection capabilities from 2D to 3D space. Despite the vast research in these two lines of research, the intersection of these two areas — open-vocabulary monocular 3D detection, referred to as OVMono3D (Fig. 1d) — remains largely unexplored.
In this work, we aim to fill this gap. The OVMono3D task involves detecting and localizing objects of any categories in the metric 3D space, including novel categories unseen during training. We identify two primary challenges in this task: first, the lack of a well-defined task setting and standardized evaluation protocol, which complicates consistent model assessment; and second, the scarcity and limited quality of labeled 3D data, as annotating detailed 3D bounding boxes is labor-intensive (Fig. 2 and Fig. 3).
To address these challenges, we first propose a formal definition of the task and a simple protocol to mitigate the impact of annotation issues (e.g., missing labels) in evaluations (Sec. 3). Our evaluation metric enables direct performance assessment without exhaustive 3D annotations, which is especially useful for efficient real-world evaluation of a detection system, where annotating every object is impractical and error-prone.
One unique challenge in OVMono3D is the lack of large-scale 3D datasets with high-quality annotations (Fig. 2). For this reason, we adopt an approach that decouples the tasks of recognizing and localizing objects in 2D from estimating their 3D bounding boxes. We first consider a simple training-free method which un-projects 2D detections from off-the-shelf open-vocabulary 2D detectors [12, 32] into 3D via geometric principles, referred to as OVMono3D-GEO (Fig. 4a). Despite of its simplicity and effectiveness in general cases, we find this approach lacks of robustness to object occlusion.
To tackle this issue, we then introduce a data-driven approach, OVMono3D-LIFT, which learns to lift 2D detections to 3D (Fig. 4b). We comprehensively explore its design space, examining choices such as network backbones, 2D base detectors, and dataset size. Our experiments show that OVMono3D-LIFT achieves state-of-the-art performance in detecting novel objects in 3D, and demonstrates strong generalization on in-the-wild images (Fig. 7).
We summarize our contributions as follows:
We conduct the first comprehensive study on open-vocabulary monocular 3D object detection, formally define the task and evaluation metrics, and extensively explore the design space of several baseline methods.
The proposed OVMono3D-LIFT effectively decouples the tasks of 2D recognition and localization from 3D spatial estimation, achieving superior performance compared to alternative baseline methods.
We provide an in-depth analysis of design choices and demonstrate the effectiveness of our approach in zero-shot monocular 3D detection on in-the-wild images.
2 Related Work
Open-Vocabulary 2D Object Detection.
The goal of this task is to recognize and localize objects in 2D images beyond a fixed set of predefined categories. Leveraging large-scale 2D datasets [49, 25, 39], this field has seen considerable advancement. Some approaches [57, 18, 36, 35] employ pre-trained vision-language models [24, 41], using frozen text features to detect novel categories. Other methods [28, 60, 32, 12] are pre-trained on extensive detection, grounding, and caption data to align region-text features. For example, Grounding DINO [32] incorporates grounded pre-training with cross-modal feature fusion, while YOLO-World [12] uses region-text contrastive loss and re-parameterization to enhance both accuracy and efficiency. However, such large-scale annotations are costly in 3D detection; thus, we investigate methods to adapt existing open-vocabulary 2D detectors for 3D detection.
Open-Vocabulary 3D Object Detection.
This task seeks to identify objects from any category in 3D, including those unseen during training. Prior research [33, 34, 7, 8, 58, 70, 63, 61, 40] primarily focuses on 3D detection tasks with 3D point clouds as input. [33] first proposes an open-vocabulary 3D detector using image-level class supervision from ImageNet1K [48]. [34] utilizes 2D bounding boxes from a pre-trained 2D detector [68] to build pseudo-3D label for object localization. [58] leverages various 2D foundation models to enhance the performance of 3D open-vocabulary detection. More recently, OV-Uni3DETR [54] proposes a multi-modal open-vocabulary 3D detector that accommodates both point clouds and images as the input. In contrast, our work focuses on the monocular 3D detection task that only requires RGB images as input and does not assume the availability of point cloud data at training or inference phases.
Monocular 3D Object Detection
refers to the task of identifying and localizing objects within a scene using 3D bounding boxes derived from single-view images. Early research in this domain primarily targeted either outdoor [11, 66, 67, 52, 10, 64, 22, 53] or indoor [13, 23, 37, 46] environments, focusing on specialized applications such as autonomous driving in urban areas and room layout estimation. The release of the extensive Omni3D dataset enabled Cube R-CNN [5] to pioneer unified monocular 3D object detection, allowing detectors to be effectively applied across multiple scene types and significantly enhancing their generalization capabilities. Following this, UniMODE [30] introduced the first successful BEV-based monocular 3D object detector capable of operating in both indoor and outdoor settings, demonstrating increased versatility in diverse environments. However, most of these approaches suffer from limited generalizability, and even advanced models like Cube R-CNN [5] and UniMODE [30] are constrained by a closed vocabulary, restricting their ability to recognize or detect object classes that were not included during training. To address this limitation, our work focuses on exploring the potential of monocular open-vocabulary 3D detection.
3 Open Vocabulary Monocular 3D Detection
3.1 Task Definition
Formally, given an input image and a text prompt consisting of all possible category names from an open vocabulary , the objective of this task is to predict a set of objects , where:
is the predicted category label for object .
is the parameters of the 3D bounding box in the camera coordinate system, defined in metric space.
The vocabulary is divided into two subsets: base categories , which are seen during training, and novel categories , which are unseen during training and only evaluated during inference. The model is trained to predict the 3D bounding boxes for objects in and generalize to without additional fine-tuning.
3.2 Evaluation Metrics
Standard Metrics.
Mean Average Precision (AP) based on Intersection over Union (IoU) is the widely used standard metric for closed-vocabulary 3D object detection tasks [5, 30]. However, directly applying traditional evaluation protocols to OVMono3D presents challenges due to common issues in 3D dataset annotations:
Naming ambiguity: Objects may be labeled with different conventions or policies during annotation (e.g., table vs. desk). Standard open-vocabulary 2D detection methods typically prompt 2D detectors with exhaustive lists of possible categories, which can lead to correct predictions with class names that do not align with the dataset annotations, especially for vaguely defined or overlapping categories, as shown in Fig. 3b.
Target-Aware Metrics.
To address these issues, we propose a customized evaluation approach that considers only the categories with ground-truth labeled instances in each image. Specifically, instead of providing the 2D detector with an exhaustive list of possible categories, we only prompt it with category names that exist in the annotations for each image. Since human annotators typically label all instances of the same category within a single image, categories with ground-truth annotations are likely to be fully annotated.
Our target-aware evaluation prevents the negative effects of missing annotations on metrics by focusing only on categories present in the ground-truth annotations. It also excludes irrelevant categories that could distract the 2D detector. As shown in Fig. 3c, by concentrating on relevant categories, this approach mitigates the impact of missing annotations and reduces category ambiguity, enabling a more accurate assessment of the model’s 3D detection capabilities.
4 Methodology
Our approach builds on two state-of-the-art frameworks: Cube R-CNN [5] for monocular 3D detection and Grounding DINO [32] for open-vocabulary (OV) 2D detection. In this section, we first provide a brief overview of these two frameworks (Sec. 4.1); we then introduce our two proposed methods for addressing OVMono3D (Sec. 4.2 - 4.3).
4.1 Preliminaries
Cube R-CNN [5] is a state-of-the-art monocular 3D detection model, which is trained from a large-scale 3D dataset (i.e., Omni3D). It extends Faster R-CNN [44] with a 3D cube head. Using 2D region proposals as input, the cube head employs ROI poolers to extract local image features and then predicts 3D bounding boxes with MLPs. Each 3D bounding box is defined by its geometric parameters:
Position: , the 3D coordinates of the object’s center in metric space.
Dimensions: , representing width, height, and length in meters.
Orientation: Represented by a continuous 6D allocentric rotation [69].
The training objective of Cube R-CNN is defined as:
where includes the classification and bounding box regression losses from the 2D detection head [44], is the loss from the 3D cube head, and denotes the uncertainty score. The 3D loss consists of disentangled losses for each 3D attribute [50]:
where represents variable groups for the 2D center shift, depth, dimensions, and rotation. Each disentangled loss isolates the error of a specific group by substituting other predicted variables with ground-truth values when constructing the predicted 3D bounding box . The holistic loss compares the predicted 3D bounding box with the ground truth using the Chamfer Loss:
Grounding DINO [32] is a leading framework for OV 2D detection, which combines the Transformer-based DINO detector [59] with grounded pretraining. We adopt a pretrained Grounding DINO as our default OV 2D detection model due to its superior performance and strong zero-shot generalization capabilities. An ablation study on the effect of different base 2D detection models on our approach is provided in Sec. 5.
4.2 OVMono3D-GEO: Geometric 2D Unprojection
To perform OV 3D detection from a single image, a simple approach is to unproject 2D detections into 3D using geometric principles. Specifically, given an input image , a text prompt , and 2D bounding boxes with category labels predicted by an OV 2D detector [32], the method proceeds as follows.
For each detected object , an instance segmentation mask is obtained using a segmentation model (e.g., SAM [26]), and a metric depth estimation model (e.g., Depth Pro [4]) generates the depth map . Pixels within are unprojected into 3D space using the camera intrinsic matrix to form a point cloud . Here, and denote the pixel coordinates in the image, and each 3D point is computed as:
To estimate the 3D bounding box parameters , Principal Component Analysis (PCA) determines the orientation , dimensions , and centroid of . DBSCAN [14] removes outliers, improving robustness against noisy masks and depth predictions.
This geometric approach reconstructs 3D bounding boxes from 2D detections without relying on learned 3D attributes, providing a non-learning-based benchmark for evaluating OV 3D object detection models. Its accuracy depends on the quality of and , and is sensitive to occlusions, which may cause incomplete point clouds and inaccuracies in the reconstructed bounding boxes.
4.3 OVMono3D-LIFT: Class-Agnostic 2D Lifting
Given the limitations of geometric-based methods, we propose leveraging monocular 3D annotations to build data-driven models. As depicted in Fig. 2, 3D annotations are far fewer than 2D annotations. To address this, we utilize off-the-shelf 2D data-driven priors to enhance 3D performance. Specifically, we introduce OVMono3D-LIFT, which decouples OV monocular 3D detection into two stages: (1) recognizing and localizing objects in 2D using off-the-shelf OV 2D detectors, and (2) class-agnostically lifting the 2D bounding boxes to 3D cuboids.
Given an image , a text prompt , and 2D bounding boxes with category labels predicted by an OV 2D detector [32], OVMono3D-LIFT employs a pretrained Vision Transformer (ViT-B/14) backbone to extract an image feature map , where is the feature dimension. A Simple Feature Pyramid (SFP) [29] module generates three feature maps with scaling factors 0.5, 1, and 2, forming a multi-scale hierarchical pyramid. The predicted 2D bounding boxes are used as regions of interest for ROI pooling, extracting local object features of size . These features are fed into a cube head, as described in Sec. 4.1, to predict 3D attributes. Unlike [5], our attribute prediction head is class-agnostic, while their method employs class-specific layers and per-class average size statistics, which limit generalization to open vocabularies. The training objective is identical to Cube R-CNN, as described in Sec. 4.1.
Baseline Methods.
An alternative data-driven approach is to directly extend OV 2D detectors to 3D. Specifically, we use the output features from the final decoder layer of Grounding DINO [32] as object representations, which are then processed by an MLP to predict 3D attributes. While straightforward, this method does not leverage pretrained image encoders (e.g. DINOv2 [38]). Therefore, we use it as a baseline for comparison.
We evaluate two variants of this method: (1) G-DINO-3D (FZ), where only the additional MLP is trainable and other parameters are frozen, and (2) G-DINO-3D (FT), where we fine-tune all network parameters. Both variants are initialized with the pretrained weights of Grounding DINO and follow the same training objective as Cube R-CNN, as outlined in Sec. 4.1.
5 Experiments
5.1 Experimental Setup
Datasets.
The experiments are conducted using the Omni3D [5] dataset, the largest publicly available dataset for monocular 3D object detection across both indoor and outdoor scenes. Omni3D is repurposed and combined from six established sources including SUN RGB-D [51], ARKitScenes [2], Objectron [1], Hypersim [45], nuScenes [6], and KITTI [15]. It comprises a substantial 234k images, 3 million labeled 3D bounding boxes, and covers 98 distinct object categories.
During training, we use the 50 categories Cube R-CNN [5] trained on as the base categories for model training. For evaluation, we select 22 categories from the remaining classes as novel categories. These categories are chosen based on two criteria: the number of test instances and the precision of category naming. To facilitate a detailed assessment of zero-shot generalization capabilities, we further divide these categories into easy and hard subsets according to object visibility. The easy categories include board, printer, painting, microwave, tray, podium, cart, and tram, while the remaining categories constitute the hard subset.
| Methods | AP↑ | AP↑ | AP↑ | AP↑ | AP↑ | AP↑ | AP↑ |
| G-DINO-3D (FT) | 6.77 | 4.43 | 6.21 | 4.92 | 1.12 | 4.77 | 4.24 |
| G-DINO-3D (FZ) | 19.22 | 3.54 | 5.74 | 2.71 | 0.07 | 4.59 | 2.96 |
| OVMono3D-GEO | 18.72 | 9.03 | 14.88 | 7.60 | 0.43 | 12.55 | 7.01 |
| OVMono3D-LIFT | 21.59 | 16.05 | 24.27 | 14.97 | 2.40 | 20.58 | 13.47 |
Evaluations.
We report a mean AP across all 50 base categories and 22 novel categories in Omni3D and over a range of IoU3D thresholds . Unless explicitly stated otherwise, all the evaluations are under our target-aware evaluation protocols (Sec. 3.2).
Besides the standard IoU metrics, we also report the Normalized Hungarian Distance (NHD) [3] which provides a more precise evaluation. NHD computes the normalized distance between predicted and ground-truth cube corners using the Hungarian matching algorithm. NHD is computed for predictions that have a high IoU with the ground-truth 2D bounding boxes. Therefore, NHD values for different methods with varying 2D inputs are not directly comparable. To disentangle the error contributions of different predicted attributes to the overall 3D bounding box prediction, we report disentangled NHD for the position (XY), depth, dimensions, and pose. For each attribute, we compute the NHD of a disentangled 3D bounding box that uses the predicted value for that attribute while setting the other attributes to their ground-truth values, comparing it with the ground-truth cube.
Implementation Details.
Our implementation is based on Pytorch3D [43] and Detectron2 [56]. We utilize DINOv2 [38] as the image feature encoder, freezing all parameters during training. All experiments are conducted on four NVIDIA A100 GPUs. The models are trained for 116k steps with a batch size of 64. We employ SGD with an initial learning rate of 0.012, which decays by a factor of 10 at 60% and 80% of the training duration. During training, image augmentations including random horizontal flipping and resizing are applied.
| AP↑ | AP↑ | AP↑ | AP↑ | AP↑ | |
| Cube R-CNN [5] | 32.31 | 26.78 | 35.71 | 28.72 | 10.48 |
| OVMono3D-LIFT | |||||
| + Grounding DINO [32] | 21.96 | 24.77 | 33.23 | 26.33 | 9.38 |
| + Cube R-CNN [5] | 32.45 | 26.44 | 35.30 | 28.40 | 10.26 |
| + Ground Truth | 100 | 29.07 | 39.74 | 30.38 | 10.39 |
5.2 Model Performance
Fig. 5 shows qualitative results on the Omni3D test set. In comparison with Cube R-CNN [5], OVMono3D-LIFT detects not only objects of base categories, but also novel categories that are unseen at the training time.
Novel Category Performance.
Tab. 1 shows that OVMono3D-LIFT outperforms the baseline methods on novel categories by a large margin. OVMono3D-GEO demonstrates lower performance due to its sensitivity to occlusions and inaccuracies in metric depth map estimation. G-DINO-3D (FZ) freezes the 2D detection parameters, resulting in object features that are not well-suited for 3D detection. Conversely, G-DINO-3D (FT) fine-tunes the model, degrading its original OV detection capabilities, since monocular 3D detection lacks the extensive detection and grounding annotations available in 2D. Therefore, these approaches are suboptimal. In contrast, OVMono3D-LIFT effectively leverages 2D data-driven priors, including OV 2D detectors and self-supervised 2D features, by decoupling the OVMono3D task into OV 2D detection and 2D-to-3D unprojection. This approach mitigates issues related to low-quality and scarce annotations in 3D, leading to superior performance.
Base Category Performance.
Tab. 2 compares OVMono3D-LIFT with Cube R-CNN on base categories using different 2D bounding box inputs. When using the OV 2D detector’s predictions as input, OVMono3D-LIFT shows lower performance due to the OV 2D detector’s lower zero-shot performance on the Omni3D test set. In contrast, when using Cube R-CNN’s 2D predictions, OVMono3D-LIFT achieves comparable performance to Cube R-CNN. Notably, Cube R-CNN utilizes per-category priors for object dimensions and depth, highlighting the superior performance of OVMono3D-LIFT without relying on such priors. Additionally, we provide results with ground-truth 2D inputs as a reference.
| Original Evaluation | Target-Aware Evaluation | |||||||
| Models | AP | AP | AR | AR | ||||
| Cube R-CNN (Base) | 27.65 | 23.68 | 42.35 | 37.56 | 32.31+4.66 | 26.78+3.10 | 42.35+0.00 | 37.56+0.00 |
| OVMono3D-LIFT (Base) | 12.27 | 17.22 | 27.19 | 33.45 | 21.96+9.69 | 24.77+7.55 | 40.13+12.94 | 47.28+13.83 |
| OVMono3D-LIFT (Novel) | 7.01 | 4.22 | 17.86 | 15.58 | 21.59+14.58 | 16.05+11.83 | 37.88+20.02 | 36.85+21.27 |
| Methods | NHD | XY | Depth | Size | Pose |
| OVMono3D-GEO | 5.47 | 2.23 | 4.30 | 2.34 | 2.04 |
| OVMono3D-LIFT | 5.72 | 2.28 | 4.73 | 1.05 | 0.55 |
| 2D Box Input | AP↑ | AP↑ | AP↑ | AP↑ | AP↑ |
| Ground Truth | 100 | 17.21 | 25.71 | 16.30 | 2.79 |
| YOLO-World | 20.14 | 12.82 | 20.25 | 11.91 | 1.63 |
| Grounding DINO | 21.59 | 16.05 | 24.27 | 14.97 | 2.40 |
5.3 Analysis
Evaluation Protocol.
Tab. 3 shows that our target-aware evaluation leads to increased metrics on both base and novel category settings. Even for the original Cube R-CNN, our method demonstrates improvements in both AP and AP, indicating its effectiveness in eliminating the adverse effects of missing annotations. On novel categories, our target-aware evaluation shows considerable metric value gains. This improvement is particularly important because novel categories are often infrequent and smaller in size, making them harder for the model to detect when prompted with an extensive list of possible categories.
Disentangled Metrics.
Tab. 4 exhibits the overall and disentangled NHD for different attribute predictions. For these two methods, object depth prediction consistently contributes the most error to the overall prediction. This indicates that object depth estimation is the primary bottleneck of our task. Furthermore, the size and pose predictions in OVMono3D-GEO exhibit larger errors, indicating that geometry-based methods are not effective for these predictions. This underscores the necessity of developing learning-based methods.
2D Bounding Box Input.
Tab. 5 evaluates the impact of 2D detectors on both 2D and 3D detection performance for novel categories. We test two state-of-the-art 2D detectors: YOLO-World [12] and Grounding DINO [32], alongside ground-truth 2D bounding boxes as reference. Grounding DINO consistently outperforms YOLO-World across all evaluation metrics. Notably, the 3D lifting results from Grounding DINO’s zero-shot 2D detection are only 1% lower than the performance with ground-truth 2D input in terms of AP, demonstrating that Grounding DINO is a preferable off-the-shelf 2D detector for OVMono3D.
Pre-trained Backbone Selection.
Tab. 6 shows the impact of pretrained feature extractors on 3D detection performance for novel categories. DINOv2 achieves the best performance across all evaluation metrics. This result underscores the effectiveness of DINOv2’s representations for the 3D detection task. Our findings align with recent studies that highlight DINOv2’s strong zero-shot capabilities in understanding depth, multi-view correspondences, and relative pose [27, 17, 62], demonstrating that DINOv2’s features are highly suitable for this task.
| Feature Extractor | Arch. | Supervision | Dataset | AP↑ | AP↑ | AP↑ | AP↑ |
| MAE [21] | ViT-B/16 | SSL | ImageNet1k | 7.72 | 11.95 | 6.43 | 0.80 |
| CLIP [41] | ViT-B/16 | VLM | WIT-400M | 9.02 | 14.19 | 8.44 | 0.50 |
| MiDas [42] | ViT-L/16 | Depth | MIX-6 | 10.65 | 17.42 | 9.87 | 0.69 |
| DINOv2 [38] | ViT-B/14 | SSL | LVD-142M | 16.05 | 24.27 | 14.97 | 2.40 |
Training Data Scaling Law.
Fig. 6 reports our model’s AP score as a function of the training data size. This underscores the critical importance of dataset size in open vocabulary monocular 3D object detection tasks and suggests that our model may achieve even better performance with more extensive training data.
Do synthetic data help?
Tab. 7 presents the effect of synthetic data on the performance of OVMono3D-LIFT. When synthetic data is incorporated alongside real data, a modest yet meaningful increase of 1 AP point is observed in detecting objects from seen categories, while performance on novel categories remains largely unchanged. These findings suggest that, while synthetic data can enhance model performance in closed-vocabulary 3D object detection tasks, its benefits are minimal for detecting unseen objects, thereby limiting its usefulness in open-vocabulary 3D object detection scenarios.
| Data | #Images | AP | AP |
| Synthetic | 55k | 7.14 | 7.33 |
| Real | 120k | 23.78 | 16.20 |
| Synthetic+Real | 175k | 24.77 | 16.05 |
5.4 Zero-Shot Generalization Performance
Fig. 7 presents OVMono3D-LIFT prediction on in-the-wild images from COCO [31]. We visualize with intrinsics of , , , where is the input image resolution. The results show 2D projections of our predictions are well-aligned with the objects, and their top-down views closely match the visual scene layout. Even on completely out-of-distribution categories such as elephant and apple, our method produces promising results. This suggests that our model demonstrates zero-shot generalization ability on real-world images.
6 Discussion
We investigate OVMono3D, an under-explored task of recognizing and localizing objects from any categories in 3D using a single image. We identify unique challenges, notably data scarcity and limitations of standard evaluation metrics. We propose simple yet effective approaches, including geometry-based OVMono3D-GEO and learning-based OVMono3D-LIFT — both decouple 2D detection from 3D bounding box prediction. Such design enables the full utilization of off-the-shelf open-vocabulary 2D detectors pre-trained on large-scale datasets. Our analysis pinpoints the key components in the framework, including pre-trained image encoder, base 2D detectors, and the impact of dataset scaling. We further demonstrate the zero-shot generalizability of our approach on in-the-wild images. Our findings suggest that dataset scale and accurate depth perception remain the major bottlenecks in this task. One promising direction may be to develop unsupervised learning that harness the abundance of unlabeled images. We hope this work inspires future research toward advancing this task.
Acknowledgement.
The authors acknowledge Research Computing at The University of Virginia for providing computational resources and technical support that have contributed to the results reported within this publication.
Supplementary Material
Sec. A presents per-category performance on novel classes in 2D and 3D for OVMono3D-GEO and OVMono3D-LIFT. Sec. B provides additional qualitative visualizations on various datasets and compares predictions between OVMono3D-GEO and OVMono3D-LIFT. Sec. C discusses failure cases, highlighting challenges such as occlusions, out-of-distribution objects, and small or distant instances.
Appendix A Per-category Performance on novel classes
We show per-category performance on Average Precision (AP) and Average Recall (AR) on both 2D and 3D for OVMono3D-GEO and OVMono3D-LIFT in Tabs. 8 and 9, respectively.
| Category | AP | AP | AR | AR |
| Board | 29.41 | 2.06 | 34.05 | 7.78 |
| Printer | 31.05 | 8.96 | 41.77 | 22.99 |
| Painting | 33.39 | 2.09 | 44.80 | 9.04 |
| Microwave | 31.60 | 14.06 | 39.45 | 28.27 |
| Tray | 15.65 | 3.40 | 22.30 | 13.11 |
| Podium | 46.66 | 36.71 | 53.85 | 46.92 |
| Cart | 31.49 | 32.52 | 38.25 | 41.75 |
| Tram | 4.72 | 0.61 | 6.91 | 1.58 |
| Easy Categories | 28.00 | 12.55 | 35.17 | 21.43 |
| Monitor | 18.84 | 6.55 | 30.02 | 19.36 |
| Bag | 13.98 | 8.20 | 23.46 | 21.45 |
| Dresser | 27.43 | 14.87 | 38.22 | 30.86 |
| Keyboard | 12.06 | 7.23 | 25.41 | 18.56 |
| Drawers | 19.87 | 9.29 | 32.67 | 21.42 |
| Computer | 15.73 | 5.73 | 25.00 | 19.70 |
| Kitchen Pan | 5.56 | 7.36 | 14.22 | 16.67 |
| Potted Plant | 15.05 | 10.65 | 28.06 | 27.10 |
| Tissues | 5.88 | 3.04 | 11.55 | 9.86 |
| Rack | 9.19 | 3.54 | 12.50 | 8.17 |
| Toys | 10.35 | 8.84 | 21.94 | 19.17 |
| Phone | 3.30 | 5.00 | 9.52 | 12.62 |
| Soundsystem | 13.09 | 7.43 | 20.47 | 19.53 |
| Fireplace | 17.48 | 0.42 | 39.51 | 3.86 |
| Hard Categories | 13.42 | 7.01 | 23.75 | 17.74 |
| All Categories | 18.72 | 9.03 | 27.91 | 19.08 |
| Category | AP | AP | AR | AR |
| Board | 38.90 | 4.83 | 54.51 | 20.52 |
| Printer | 31.02 | 16.23 | 44.08 | 43.95 |
| Painting | 34.99 | 2.80 | 50.23 | 19.77 |
| Microwave | 32.76 | 30.31 | 42.91 | 56.27 |
| Tray | 17.75 | 10.11 | 32.79 | 30.16 |
| Podium | 45.63 | 48.37 | 57.18 | 73.08 |
| Cart | 37.54 | 47.31 | 50.75 | 71.50 |
| Tram | 22.86 | 4.71 | 34.00 | 9.15 |
| Easy Categories | 32.68 | 20.58 | 45.81 | 40.55 |
| Monitor | 21.46 | 9.44 | 38.64 | 35.44 |
| Bag | 15.90 | 15.61 | 30.22 | 36.42 |
| Dresser | 30.54 | 29.08 | 48.49 | 60.26 |
| Keyboard | 14.60 | 9.13 | 36.71 | 25.07 |
| Drawers | 21.48 | 43.04 | 43.00 | 68.33 |
| Computer | 16.80 | 7.44 | 35.30 | 28.56 |
| Kitchen Pan | 6.81 | 9.98 | 21.57 | 35.10 |
| Potted Plant | 12.94 | 6.66 | 33.23 | 24.62 |
| Tissues | 8.81 | 12.45 | 21.55 | 32.68 |
| Rack | 13.02 | 10.21 | 27.33 | 35.33 |
| Toys | 10.91 | 5.24 | 29.31 | 22.08 |
| Phone | 6.23 | 3.89 | 22.14 | 16.43 |
| Soundsystem | 14.52 | 13.22 | 31.86 | 31.86 |
| Fireplace | 19.50 | 13.16 | 47.62 | 34.06 |
| Hard Categories | 15.25 | 13.47 | 33.36 | 34.73 |
| All Categories | 21.59 | 16.05 | 37.88 | 36.85 |
Appendix B More Qualitative Results
Additional qualitative visualizations of OVMono3D-LIFT are provided for Omni3D [5] outdoor, indoor subsets, and COCO [31] in-the-wild images in Figs. 8, 9 and 10, respectively.
Fig. 11 illustrates a comparison between the predictions of OVMono3D-GEO and OVMono3D-LIFT. OVMono3D-GEO derives object depth from an estimated metric depth map, yielding better relative depth consistency with scene layout (e.g., Fig. 11c). However, it estimates dimensions and poses based on visible object parts, leading to biases. For instance, it struggles with occlusions (e.g., the door in Fig. 11d), limited surface visibility (e.g., ovens in Fig. 11e), and noisy depth maps (e.g., the farthest chair in Fig. 11f). In contrast, OVMono3D-LIFT, leveraging learned priors, is more robust in such scenarios. Future work could integrate these methods to mitigate their respective limitations.
Appendix C Failure Cases
Fig. 12 shows failure cases of OVMono3D-LIFT on COCO [31] images. In Fig. 12a, the relative position from a top-down view is incorrect, indicating that our model sometimes predicts the wrong object depth. In Fig. 12b, our model fails to predict the correct size and pose for the bear, suggesting that it may struggle with totally out-of-distribution objects. In Fig. 12c, our model fails to detect the person and bus in the distance, indicating that it may not perform well on small and distant objects. In Fig. 12d, our method fails to identify the mirror and incorrectly detects the object in the mirror. These failure cases suggest that our model still has room for improvement. Future research could explore better model architectures and weakly supervised learning techniques to address these shortcomings.